OpenStudy (anonymous):

P=square root of (100-2Q) , isolate the Q

OpenStudy (anonymous):

\[P^2=\sqrt{100-2Q}^2\]\[P^2-100=-2Q\]\[Q=-\frac{100-P^2}{2}\]

OpenStudy (anonymous):

last line should be \[Q = \frac{100-P^2}{2}\]

OpenStudy (anonymous):

thanks